what is the ordered pair for the equation y=8x-4, x=4

Answer by nerdybill(7384)

what is the ordered pair for the equation y=8x-4, x=4
.
They give us the x-value. All we need to do is find the y-value. Do this by plugging in the x-value to:
y=8x-4
y=8(4)-4
y=32-4
y=28
.
answer:
(4, 28)